Hello! doing a calculation but stumped at the moment. If $3,325.00 represents .35 (35%) of the total cost, what is the TOTAL COST?
Thanks for your help!
AD

Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

If $3,325.00 represents .35 (35%) of the total cost, what is the TOTAL COST?
------
Let the total cost be "x":
-----
0.35*x = 3325
x = 3325/0.35 = $9500.00
---------------
Cheers,
Stan H.
------------------

Answer by ikleyn(52782)